OpenStudy (anonymous):

Which of the following are arithmetic sequences? a. 1, 2, 3, 4, 5, 6, ... b. 2, 2, 2, 2, 2, ... c. 2, 6, 18, 54, ... d. 100, 230, 360, 490, ... *** check ALL that apply

OpenStudy (anonymous):

first

OpenStudy (anonymous):

Arithmetic means that the pattern is the same yes?

OpenStudy (anonymous):

second

OpenStudy (anonymous):

IF so, I have A and C so far cuz A is adding one each time and C is x3 right???

OpenStudy (anonymous):

Idk is B an answer though?? I got confused on that because there is no change... idk could u pls explain ? :)

jimthompson5910 (jim_thompson5910):

2 = 2+0 2 = 2+0 ... etc So the common difference is 0

jimthompson5910 (jim_thompson5910):

So 2, 2, 2, 2, 2, ... is an arithmetic sequence

OpenStudy (anonymous):

And I have no idea about D.... But am i right that arithmetic means constant difference?

OpenStudy (anonymous):

oh so B is an answer as well? :) Because +0 can be a difference?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

Oh cool! So far we have A,B, and C???? is that right?

OpenStudy (anonymous):

arithmetic means adding and subtracting

OpenStudy (anonymous):

c cant work

OpenStudy (anonymous):

ohhhhh i seee (thanks!)

jimthompson5910 (jim_thompson5910):

C is NOT an arithmetic sequence

jimthompson5910 (jim_thompson5910):

6 - 2 = 4 but 18 - 6 = 12 so there is no common difference

OpenStudy (anonymous):

ohhhh how come it's not?? is it cuz like you said it's only adding and subtraction?? so no multiplication?

OpenStudy (anonymous):

cuz i got multiply by 3 as the common thingy...

jimthompson5910 (jim_thompson5910):

that's the common ratio

jimthompson5910 (jim_thompson5910):

so it means that C is a geometric sequence instead

OpenStudy (anonymous):

ohhhh geometric means multiplication?? DOes that also include division?

OpenStudy (anonymous):

And what about D? I couldn't find anything for D.... ;(

jimthompson5910 (jim_thompson5910):

yes, multiplication and division are linked (since division is really multiplication of the multiplicative inverse)

OpenStudy (anonymous):

oh i see now.... oh wait nvm... is D an arithmetic sequence cuz it's +130??? each time???

jimthompson5910 (jim_thompson5910):

230 - 100 = ? 360 - 230 = ? 390 - 360 = ?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

So the answer to this question is A,B, and D?

jimthompson5910 (jim_thompson5910):

oops, 490 - 360 = ?

jimthompson5910 (jim_thompson5910):

yes, that's correct

OpenStudy (anonymous):

oh okay!! I get it now!! :) Thanks!! :) Can i message u the next link when i post it in a bit? :)

OpenStudy (anonymous):

oh and can u pls give @Jonask a medal for me @jim_thompson5910 ??? cuz i can only give one :(